# Least-Squares Neural Network (LSNN) Method for Linear Advection-Reaction Equation:

### v2

@@ 4.2 A problem with a piecewise smooth solution
$\Gamma_{-}$ $=$ $\{(x,0):x\in(0,1)\}$
and $\mbox{and }\,\,g(x,y)$ $=$ $\left\{\begin{array}[]{rl}0,&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1/2)\},\\[5.69054pt] 2,&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
respectively. The exact solution of this test problem is
$u(x,y)=\left\{\begin{array}[]{rl}1-e^{-y},&(x,y)\in\Omega_{1},\\[5.69054pt] 1+e^{-y},&(x,y)\in\Omega_{2}.\end{array}\right.$ (19)
50000 iterations were implemented with 2–20–20–1 ReLU NN functions. The numerical results are presented in Figs. 3 and 2. Unlike the previous test problem, the exact solution (Fig. 3(c)) consists of two non-constant smooth parts. The LSNN method is capable of approximating the solution accurately without oscillation (Figs. 3(b), 3(c), 3(d), and 2). The 3-layer ReLU NN function approximation has a partition (Fig. 3(e)) similar to that in Section 4.1 with the second-layer breaking hyperplanes on both sides for approximating the two non-constant smooth parts of the solution.
@@ 4.3 A problem with a piecewise smooth inflow boundary
$\Gamma_{-}$ $=$ $\{(x,0):x\in(0,1)\}$
and $\mbox{and }\,\,g(x,y)$ $=$ $\left\{\begin{array}[]{rl}1-\sin(2\pi x),&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1/2)\},\\[5.69054pt] 5/2-x,&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
respectively. The exact solution of this test problem is
$u(x,y)=\left\{\begin{array}[]{rl}1-\sin(2\pi x)e^{-y},&(x,y)\in\Omega_{1},\\[5.69054pt] 1+(3/2-x)e^{-y},&(x,y)\in\Omega_{2}.\end{array}\right.$ (21)
100000 iterations were implemented with 2–40–40–1 ReLU NN functions. The numerical results are presented in Figs. 4 and 3. Since the solution on the inflow boundary consists of two non-constant smooth curves, we increased the number of hidden neurons to obtain a more accurate solution. Figs. 4(c), 4(d), and 3 show that the approximation is accurate pointwise and in average. The traces (Fig. 4(b)) on $y=0.5$ exhibit no oscillation and we note a few corners on the curve, verifying that the ReLU NN function approximation is a CPWL function. The partition generated by the breaking hyperplanes (Fig. 4(e)) of the approximation shows how the exact solution was approximated.
@@ 4.4 A problem with a piecewise constant advection velocity field
The advective velocity field is a piecewise constant field given by
$\bm{\beta}(x,y)=\left\{\begin{array}[]{rl}(-1,\sqrt{2}-1)^{T},&(x,y)\in\Upsilon_{1},\\[5.69054pt] (1-\sqrt{2},1)^{T},&(x,y)\in\Upsilon_{2}.\end{array}\right.$ (22)
The inflow boundary and the inflow boundary condition are given by
$\Gamma_{-}$ $=$ $\{(1,y):y\in(0,1)\}\cup\{(x,0):x\in(0,1)\}$
and $\mbox{and }\,\,g(x,y)$ $=$ $\left\{\begin{array}[]{rl}x\exp(x/(\sqrt{2}-1)),&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1)\},\\[5.69054pt] (11+(\sqrt{2}-1)y)\exp(1/(\sqrt{2}-1)),&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
respectively. Let
@@
The exact solution of this test problem is
$u(x,y)=\left\{\begin{array}[]{rl}(y+(\sqrt{2}-1)x)\exp(\sqrt{2}x+y),&(x,y)\in\widehat{\Upsilon}_{11},\\[5.69054pt] (y+(\sqrt{2}-1)x+10)\exp(\sqrt{2}x+y),&(x,y)\in\widehat{\Upsilon}_{12},\\[5.69054pt] (x+(\sqrt{2}-1)y)\exp(x/(\sqrt{2}-1)),&(x,y)\in\widehat{\Upsilon}_{21},\\[5.69054pt] (x+(\sqrt{2}-1)y+10)\exp(x/(\sqrt{2}-1)),&(x,y)\in\widehat{\Upsilon}_{22}.\end{array}\right.$ (24)
300000 iterations were implemented with 2–450–1 and 2–40–40–1 ReLU NN functions. The numerical results are presented in Figs. 5 and 4. This example compares the approximation differences with the one-hidden-layer NN (a universal approximator) with the same number of degrees of freedom. The exact solution (Fig. 5(b)) consists of four non-constant smooth parts and has a non-constant jump along two connected line segments (Fig. 5(a)). The traces (Fig. 5(f)) of the exact and numerical solutions, the approximation (Fig. 5(d)), and Table 4 indicate that the 3-layer ReLU NN function approximation is accurate pointwise and in average. Most of the second-layer breaking hyperplanes (Fig. 5(h)) are along the discontinuous interface, which correspond to the sharp transition layer of the approximation for the jump. On the other hand, Figs. 5(c), 5(e), 5(g), and 4 show that the one-hidden-layer NN failed to approximate the solution, especially around the interface.
@@ 4.5 A problem with a variable advection velocity field
The advective velocity field is a variable field given by
$\bm{\beta}(x,y)=(1,2x),\,\,(x,y)\in\Omega.$ (25)
The inflow boundary and the inflow boundary condition are given by
$\Gamma_{-}$ $=$ $\{(0,y):y\in(0,1)\}\cup\{(x,0):x\in(0,1)\}$
and $\mbox{and }\,\,g(x,y)$ $=$ $\left\{\begin{array}[]{rl}y+2,&(x,y)\in\Gamma^{1}_{-}\equiv\{(0,y):y\in[\frac{1}{5},1)\},\\[5.69054pt] (y-x^{2})e^{-x},&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
respectively. The exact solution of this test problem is
$u(x,y)=\left\{\begin{array}[]{rl}(y-x^{2})e^{-x},&(x,y)\in\Omega_{1}\equiv\{(x,y)\in\Omega:y<x^{2}+\frac{1}{5}\},\\[5.69054pt] (y-x^{2}+2)e^{-x},&(x,y)\in\Omega_{2}=\Omega\setminus\Omega_{1}.\end{array}\right.$ (27)
300000 iterations were implemented with 2–60–60–1 ReLU NN functions. The numerical results are presented in Figs. 6 and 5. We increased the number of hidden neurons and $\rho$ was set to $h/15$ in the finite difference quotient in Eq. 14 because of the jump along the curved interface (Fig. 6(a)). Although theoretical analysis on the convergence of the method in the case of a smooth interface was not conducted, Figs. 6(b), 6(c), 6(d), and 5 show that the LSNN method is still capable of approximating the discontinuous solution with the curved interface accurately without oscillation. Finally, again, most of the second-layer breaking hyperplanes are along the interface (Fig. 6(e)) to approximate the discontinuous jump.
@@
Now we present numerical results for a three-dimensional test problem with a constant advection velocity field whose solution is piecewise smooth along a plane segment with a non-constant jump. The test problem is defined on the domain $\Omega=(0,1)^{3}$, and approximation results are depicted on $z=0.505$.
4.6 A problem with a constant advection velocity field
The advective velocity field is a constant field given by
$\bm{\beta}(x,y,z)=(1,0,0),\,\,(x,y,z)\in\Omega.$ (28)
The inflow boundary and the inflow boundary condition are given by
$\Gamma_{-}$ $=$ $\{(0,y,z):y,z\in(0,1)\}$
and $\mbox{and }\,\,g(x,y,z)$ $=$ $\left\{\begin{array}[]{rl}1-4y,&(x,y,z)\in\Gamma^{1}_{-}\equiv\{(0,y,z)\in\Gamma_{-}:y<\frac{1}{2}\},\\[5.69054pt] 5-4y,&(x,y,z)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
respectively. The exact solution of this test problem is
$u(x,y,z)=\left\{\begin{array}[]{rl}1-4ye^{-x},&(x,y,z)\in\Omega_{1}\equiv\{(x,y,z)\in\Omega:y<\frac{1}{2}\},\\[5.69054pt] 1+(4-4y)e^{-x},&(x,y,z)\in\Omega_{2}=\Omega\setminus\Omega_{1}.\end{array}\right.$ (30)
100000 iterations were implemented with 3–30–30–1 ReLU NN functions (depth $\lceil\log_{2}(d+1)\rceil+1=3$ for $d=3$). The numerical results are presented in Figs. 7 and 6. The exact solution (Fig. 7(c)) is discontinuous along the plane segment $y=0.5$ (Fig. 7(a)), and was approximated accurately by the 3-layer NN (Figs. 7(b), 7(d), and 6). The behavior of the breaking hyperplanes (Fig. 7(e)) is similar to those of the previous examples around the discontinuous interface and on the subdomains.
@@ 5 Conclusion
In this paper, we used the least-squares ReLU neural network (LSNN) method for solving linear advection-reaction equation with discontinuous solution having non-constant jumps. The method, being mesh-free, requires no mesh to resolve the interfacial discontinuity. We proved theoretically that ReLU neural network (NN) functions with $\lceil\log_{2}(d+1)\rceil+1$-layer representations are capable of approximating solutions with non-constant jumps along discontinuous interfaces that are not necessarily straight lines. Our theoretical findings were validated by multiple numerical examples with $d=2,3$ and various non-constant jumps and interface shapes, demonstrating accurate performance of the method.
The approximation of discontinuous functions by NNs is also encountered in classification tasks, and our results suggest that we can achieve accurate predictions with properly designed neural network architectures. However, in this paper, we mainly focused on the depth of NNs. The approximation of discontinuous classification functions by NNs with fixed depth and width will be addressed in a forthcoming paper.
@@
$\chi({\bf x})=\chi_{0}({\bf x})+\chi_{1}({\bf x})\text{ on }\widehat{\Upsilon},$ (31)
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|b-p_{1}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}=\left(\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}}+\|b_{{\bm{\beta}}}-p_{1{\bm{\beta}}}\|^{2}_{0,\widehat{\Upsilon}}\right)^{1/2}\leq\sqrt{\frac{\varepsilon_{1}^{3}}{24}+\frac{B^{2}}{4}}\,|d|\,\sqrt{\varepsilon_{1}\varepsilon_{2}},$ (33)
@@
$\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}}=\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}_{1,\varepsilon_{2}}}+\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}_{2,\varepsilon_{2}}}=\frac{d^{2}}{24}\varepsilon_{1}^{4}\varepsilon_{2},$ (34)
@@
$p_{0}(\mathbf{x})=\frac{1}{2\varepsilon_{2}}\left(\sigma(x-x_{0}+\varepsilon_{2})-\sigma(x-x_{0}-\varepsilon_{2})\right).$ (36)
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-(p_{0}+p_{1})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq C_{1}\sqrt{\varepsilon_{1}\varepsilon_{2}}+C_{2}\sqrt{\varepsilon_{1}},$ (37)
@@
$\sup\{(u_{1}(\mathbf{x})-u_{2}(\mathbf{x})-1-b(\mathbf{x}))^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}\}x_{0}\\ +\sup\{2\gamma(\mathbf{x})^{2}(u_{2}(\mathbf{x})-u_{1}(\mathbf{x}))^{2}+2(dB)^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}\}x_{0}.$ (38)
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-(p_{0}+p_{1})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}={\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{0}+\chi_{1}-(p_{0}+p_{1})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{0}-p_{0}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}+{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{1}-p_{1}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}.$ (39)
@@
$\begin{split}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{0}-p_{0}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}&=\left(\|\chi_{0}-p_{0}\|^{2}_{0,\widehat{\Upsilon}}+\|\chi_{0\bm{\beta}}-p_{0\bm{\beta}}\|^{2}_{0,\widehat{\Upsilon}}\right)^{1/2}\\ &=\|\chi_{0}-p_{0}\|_{0,\widehat{\Upsilon}}\\ &=\frac{1}{\sqrt{6}}\,\sqrt{\varepsilon_{1}\varepsilon_{2}}.\end{split}$ (40)
@@
$\sup\{(u_{1}({\bf x})-u_{2}({\bf x})-\chi(0,j/m)-b_{j}(\mathbf{x}))^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}_{j}\},\\ \sup\{2\gamma(\mathbf{x})^{2}(u_{2}(\mathbf{x})-u_{1}(\mathbf{x}))^{2}+2(d_{j}B)^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}_{j}\}<\varepsilon_{3},$ (41)
@@
$\begin{split}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-p_{i}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}&=\left(\sum_{j=0}^{m-1}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-(p_{0j}+p_{1j})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}_{\bm{\beta}}\right)^{1/2}\\ &\leq\left(\sum_{j=0}^{m-1}\left(D_{1j}\sqrt{\varepsilon_{1}\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{1}\varepsilon_{3}}\right)^{2}\right)^{1/2}\\ &\leq\left(m\max_{0\leq j\leq m-1}\left(D_{1j}\sqrt{\varepsilon_{1}\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{1}\varepsilon_{3}}\right)^{2}\right)^{1/2}\\ &=\sqrt{m}\max_{0\leq j\leq m-1}\left(D_{1j}\sqrt{\varepsilon_{1}\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{1}\varepsilon_{3}}\right)\\ &=\max_{0\leq j\leq m-1}\left(D_{1j}\sqrt{\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{3}}\right),\end{split}$ (42)

### v3

@@ 4.2 A problem with a piecewise smooth solution
$g(x,y)=\left\{\begin{array}[]{rl}0,&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1/2)\},\\[5.69054pt] 2,&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1}.\end{array}\right.$
The exact solution of this test problem is
(15) $u(x,y)=\left\{\begin{array}[]{rl}1-e^{-y},&(x,y)\in\Omega_{1},\\[5.69054pt] 1+e^{-y},&(x,y)\in\Omega_{2}.\end{array}\right.$
The LSNN method was implemented with 50000 iterations for 2–20–20–1 ReLU NN functions. We report the numerical results in Figs. 3 and 2. Unlike the previous test problem, the exact solution (Fig. 3(c)) consists of two non-constant smooth parts. The LSNN method is capable of approximating the solution accurately without oscillation (Figs. 3(b), 3(c), 3(d), and 2). The 3-layer ReLU NN function approximation has a partition (Fig. 3(e)) similar to that in Section 4.1 with the second-layer breaking hyperplanes on both sides for approximating the two non-constant smooth parts of the solution.
@@ 4.3 A problem with a piecewise smooth inflow boundary
$g(x,y)=\left\{\begin{array}[]{rl}1-\sin(2\pi x),&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1/2)\},\\[5.69054pt] 5/2-x,&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1}.\end{array}\right.$
The exact solution of this test problem is
(16) $u(x,y)=\left\{\begin{array}[]{rl}1-\sin(2\pi x)e^{-y},&(x,y)\in\Omega_{1},\\[5.69054pt] 1+(3/2-x)e^{-y},&(x,y)\in\Omega_{2}.\end{array}\right.$
The LSNN method was implemented with 100000 iterations for 2–40–40–1 ReLU NN functions. We report the numerical results in Figs. 4 and 3. Since the solution on the inflow boundary consists of two non-constant smooth curves, we increased the size of the neural network to obtain a more accurate solution. Figs. 4(c), 4(d), and 3 show that the approximation is accurate pointwisely and in average. The traces (Fig. 4(b)) on $y=0.5$ exhibit no oscillation and we note a few corners on the curve, verifying that the ReLU NN function approximation is a CPWL function. The partition generated by the breaking hyperplanes (Fig. 4(e)) of the approximation shows how the exact solution was approximated.
@@ 4.4 A problem with a piecewise constant advection velocity field
This test problem has a piecewise constant advective velocity field given by
(17) $\bm{\beta}(x,y)=\left\{\begin{array}[]{rl}(-1,\sqrt{2}-1)^{T},&(x,y)\in\Upsilon_{1},\\[5.69054pt] (1-\sqrt{2},1)^{T},&(x,y)\in\Upsilon_{2},\end{array}\right.$
and the inflow boundary $\Gamma_{-}=\{(1,y):y\in(0,1)\}\cup\{(x,0):x\in(0,1)\}$. For the inflow boundary condition,
$g(x,y)=\left\{\begin{array}[]{rl}x\exp(x/(\sqrt{2}-1)),&(x,y)\in\Gamma^{1}_{-}\equiv\{(x,0):x\in(0,1)\},\\[5.69054pt] (11+(\sqrt{2}-1)y)\exp(1/(\sqrt{2}-1)),&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1},\end{array}\right.$
the exact solution of this test problem is
(18) $u(x,y)=\left\{\begin{array}[]{rl}(y+(\sqrt{2}-1)x)\exp(\sqrt{2}x+y),&(x,y)\in\widehat{\Upsilon}_{11},\\[5.69054pt] (y+(\sqrt{2}-1)x+10)\exp(\sqrt{2}x+y),&(x,y)\in\widehat{\Upsilon}_{12},\\[5.69054pt] (x+(\sqrt{2}-1)y)\exp(x/(\sqrt{2}-1)),&(x,y)\in\widehat{\Upsilon}_{21},\\[5.69054pt] (x+(\sqrt{2}-1)y+10)\exp(x/(\sqrt{2}-1)),&(x,y)\in\widehat{\Upsilon}_{22},\end{array}\right.$
where
@@
The LSNN method was implemented with 300000 iterations for 2–450–1 and 2–40–40–1 ReLU NN functions. We report the numerical results in Figs. 5 and 4. This example compares the approximation differences with the one-hidden-layer NN (a universal approximator) with the same number of degrees of freedom. The exact solution (Fig. 5(b)) consists of four non-constant smooth parts and has a non-constant jump along two connected line segments (Fig. 5(a)). The traces (Fig. 5(f)) of the exact and numerical solutions, the approximation (Fig. 5(d)), and Table 4 indicate that the 3-layer ReLU NN function approximation is accurate pointwisely and in average. Most of the second-layer breaking hyperplanes (Fig. 5(h)) are along the discontinuous interface, which correspond to the sharp transition layer of the approximation for the jump. On the other hand, Figs. 5(c), 5(e), 5(g), and 4 show that the one-hidden-layer NN failed to approximate the solution, especially around the interface.
@@ 4.5 A problem with a variable advection velocity field
This test problem has the variable advective velocity field $\bm{\beta}(x,y)=(1,2x),\,\,(x,y)\in\Omega$, and the inflow boundary of the problem is $\Gamma_{-}=\{(0,y):y\in(0,1)\}\cup\{(x,0):x\in(0,1)\}$. The inflow boundary condition is given by
$g(x,y)=\left\{\begin{array}[]{rl}y+2,&(x,y)\in\Gamma^{1}_{-}\equiv\{(0,y):y\in[\frac{1}{5},1)\},\\[5.69054pt] (y-x^{2})e^{-x},&(x,y)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1}.\end{array}\right.$
The exact solution of this test problem is
(19) $u(x,y)=\left\{\begin{array}[]{rl}(y-x^{2})e^{-x},&(x,y)\in\Omega_{1}\equiv\{(x,y)\in\Omega:y<x^{2}+\frac{1}{5}\},\\[5.69054pt] (y-x^{2}+2)e^{-x},&(x,y)\in\Omega_{2}=\Omega\setminus\Omega_{1}.\end{array}\right.$
The LSNN method was implemented with 300000 iterations for 2–60–60–1 ReLU NN functions. We report the numerical results in Figs. 6 and 5. Since the advective velocity field is a variable field, we increased the size of the neural network and $\rho=h/15$ was used to compute the finite difference quotient in Eq. 14 to take values in one subdomain of the partition. Although theoretical analysis on the convergence of the method in the case of a smooth interface was not conducted, Figs. 6(b), 6(c), 6(d), and 5 show that the LSNN method is still capable of approximating the discontinuous solution with the curved interface accurately without oscillation. Finally, again, most of the second-layer breaking hyperplanes are along the interface (Fig. 6(e)) to approximate the discontinuous jump.
@@
4.6 A problem with a constant advection velocity field ($d=3$)
The last test problem is a three-dimensional problem defined on the domain $\Omega=(0,1)^{3}$, and approximation results are depicted on $z=0.505$. The advective velocity field is the constant field $\bm{\beta}(x,y,z)=(1,0,0),\,\,(x,y,z)\in\Omega$, and the inflow boundary of the problem is $\Gamma_{-}=\{(0,y,z):y,z\in(0,1)\}$. The inflow boundary condition is given by
$g(x,y,z)=\left\{\begin{array}[]{rl}1-4y,&(x,y,z)\in\Gamma^{1}_{-}\equiv\{(0,y,z)\in\Gamma_{-}:y<\frac{1}{2}\},\\[5.69054pt] 5-4y,&(x,y,z)\in\Gamma^{2}_{-}=\Gamma_{-}\setminus\Gamma_{-}^{1}.\end{array}\right.$
The exact solution of this test problem is
(20) $u(x,y,z)=\left\{\begin{array}[]{rl}1-4ye^{-x},&(x,y,z)\in\Omega_{1}\equiv\{(x,y,z)\in\Omega:y<\frac{1}{2}\},\\[5.69054pt] 1+(4-4y)e^{-x},&(x,y,z)\in\Omega_{2}=\Omega\setminus\Omega_{1}.\end{array}\right.$
The LSNN method was implemented with 100000 iterations for 3–30–30–1 ReLU NN functions. Note that we still employ the two-hidden-layer NN since $\lceil\log_{2}(d+1)\rceil+1=3$ for $d=3$. We report the numerical results in Figs. 7 and 6. The exact solution (Fig. 7(c)) is piecewise smooth along the plane segment $y=0.5$ (Fig. 7(a)) with a non-constant jump, and was approximated accurately by the 3-layer NN (Figs. 7(b), 7(d), and 6). The behavior of the breaking hyperplanes (Fig. 7(e)) is similar to those of the previous examples around the discontinuous interface and on the subdomains.
@@ 5 Conclusion
In this paper, we used the least-squares ReLU neural network (LSNN) method for solving linear advection-reaction equation with discontinuous solution having non-constant jumps. The method, being mesh-free, requires no mesh to resolve the interfacial discontinuity. We proved theoretically that ReLU neural network (NN) functions with $\lceil\log_{2}(d+1)\rceil+1$-layer representations are capable of accurately approximating solutions with non-constant jumps along discontinuous interfaces that are not necessarily straight lines. Our theoretical findings were validated by multiple numerical examples with $d=2,3$ and various non-constant jumps and interface shapes, demonstrating accurate performance of the method.
The approximation of discontinuous functions by NNs is also encountered in classification tasks, and our results suggest that we can achieve accurate predictions with properly designed neural network architectures. However, as in [5], in this paper, we mainly focused on the depth of NNs. The approximation of discontinuous classification functions by NNs with fixed depth and width will be addressed in a forthcoming paper.
@@
(21) $\chi({\bf x})=\chi_{0}({\bf x})+\chi_{1}({\bf x})\text{ on }\widehat{\Upsilon},$
@@
(22) ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|b-p_{1}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}=\left(\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}}+\|b_{{\bm{\beta}}}-p_{1{\bm{\beta}}}\|^{2}_{0,\widehat{\Upsilon}}\right)^{1/2}\leq\sqrt{\frac{\varepsilon_{1}^{3}}{24}+\frac{B^{2}}{4}}\,|d|\,\sqrt{\varepsilon_{1}\varepsilon_{2}},$
@@
(23) $\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}}=\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}_{1,\varepsilon_{2}}}+\|b-p_{1}\|^{2}_{0,\widehat{\Upsilon}_{2,\varepsilon_{2}}}=\frac{d^{2}}{24}\varepsilon_{1}^{4}\varepsilon_{2},$
@@
(25) $p_{0}(\mathbf{x})=\frac{1}{2\varepsilon_{2}}\left(\sigma(x-x_{0}+\varepsilon_{2})-\sigma(x-x_{0}-\varepsilon_{2})\right).$
@@
(26) ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-(p_{0}+p_{1})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq C_{1}\sqrt{\varepsilon_{1}\varepsilon_{2}}+C_{2}\sqrt{\varepsilon_{1}},$
@@
(27) $\sup\{(u_{1}(\mathbf{x})-u_{2}(\mathbf{x})-1-b(\mathbf{x}))^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}\}x_{0}\\ +\sup\{2\gamma(\mathbf{x})^{2}(u_{2}(\mathbf{x})-u_{1}(\mathbf{x}))^{2}+2(dB)^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}\}x_{0}.$
@@
(28) ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-(p_{0}+p_{1})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}={\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{0}+\chi_{1}-(p_{0}+p_{1})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{0}-p_{0}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}+{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{1}-p_{1}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}.$
@@
(29) $\begin{split}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi_{0}-p_{0}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}&=\left(\|\chi_{0}-p_{0}\|^{2}_{0,\widehat{\Upsilon}}+\|\chi_{0\bm{\beta}}-p_{0\bm{\beta}}\|^{2}_{0,\widehat{\Upsilon}}\right)^{1/2}\\ &=\|\chi_{0}-p_{0}\|_{0,\widehat{\Upsilon}}\\ &=\frac{1}{\sqrt{6}}\,\sqrt{\varepsilon_{1}\varepsilon_{2}}.\end{split}$
@@
(30) $\sup\{(u_{1}({\bf x})-u_{2}({\bf x})-\chi(0,j/m)-b_{j}(\mathbf{x}))^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}_{j}\},\\ \sup\{2\gamma(\mathbf{x})^{2}(u_{2}(\mathbf{x})-u_{1}(\mathbf{x}))^{2}+2(d_{j}B)^{2}:\mathbf{x}\in\Omega_{1i}\cap\widehat{\Upsilon}_{j}\}<\varepsilon_{3},$
@@
(31) $\begin{split}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-p_{i}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\bm{\beta}}&=\left(\sum_{j=0}^{m-1}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\chi-(p_{0j}+p_{1j})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}_{\bm{\beta}}\right)^{1/2}\\ &\leq\left(\sum_{j=0}^{m-1}\left(D_{1j}\sqrt{\varepsilon_{1}\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{1}\varepsilon_{3}}\right)^{2}\right)^{1/2}\\ &\leq\left(m\max_{0\leq j\leq m-1}\left(D_{1j}\sqrt{\varepsilon_{1}\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{1}\varepsilon_{3}}\right)^{2}\right)^{1/2}\\ &=\sqrt{m}\max_{0\leq j\leq m-1}\left(D_{1j}\sqrt{\varepsilon_{1}\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{1}\varepsilon_{3}}\right)\\ &=\max_{0\leq j\leq m-1}\left(D_{1j}\sqrt{\varepsilon_{2}}+D_{2j}\sqrt{\varepsilon_{3}}\right),\end{split}$